OpenStudy (wwe123):

Relations & Functions---------------------------for x>3,f(x)=3x-2 and for -2≤x≤2,f(x)x^2-2 find f(0)+f(4)

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x^2-2& \text{if} & -2\leq x\leq 2 \\ 3x-2& \text{if} & x >3 \end{array} \right. \] ?

OpenStudy (wwe123):

yes and....

OpenStudy (anonymous):

then it is ok if you understand what it means how do you find \(f(0)\)? first you ask what catagory does 0 belong to, is \(-2\leq 0\leq 2\) or is \(0>3\) ? evidently it is the first one, which means you use the first formula, i.e. \[f(0)=0^2-2=-2\]

OpenStudy (anonymous):

repeat with 4, since \(4>3\) you would use the second formula to compute \(f(4)\) then add

OpenStudy (wwe123):

please demonstrate more

OpenStudy (anonymous):

f(0)= 0^2 - 2 = -2 f(4)= 3(4) - 2 = 10 then f(0) + f(4) = -2+ 10 = 8